# Splitting differential equations using Galois theory

### v1

@@ 4.1. Preliminaries on logarithmic-differential pullbacks
Let $L(b)=L\langle b\rangle$ be the (differential) field generated by $b$ over $L$ (the two are equal since $\log_{\delta}(b)\in F$). We have that $\log_{\delta}(a)\in L(b)$, but also $a\in\operatorname{acl}(L(b))$, which implies that $a$ is field-algebraic over $L(b)$. Thus there is $P\in L(b)[X]$ such that $P(a)=0$, we take $P$ monic and of minimal degree, so $P=X^{n}+\sum\limits_{i=0}^{n-1}c_{i}X^{i}$. Then we compute:
$0=\delta(P(a))=n\log_{\delta}(a)a^{n}+\sum\limits_{i=0}^{n-1}(\delta(c_{i})+ic_{i}\log_{\delta}(a))a^{i}$
and thus:
$0=n\log_{\delta}(a)P(a)-\delta(P(a))=\sum\limits_{i=0}^{n-1}(n\log_{\delta}(a)c_{i}-\delta(c_{i})-ic_{i}\log_{\delta}(a))a^{i}$
As we picked $n$ minimal, we obtain, for all $i\leq n-1$:
$n\log_{\delta}(a)c_{i}-\delta(c_{i})-ic_{i}\log_{\delta}(a)=0$
and thus, for any $c_{i}\neq 0$:
$\log_{\delta}\left(a^{n-i}\right)=\log_{\delta}(c_{i})\text{ .}$
If there is some $c_{i}\neq 0$, we obtain that $\frac{a^{n-i}}{c_{i}}\in\mathcal{C}\cap L(b)(a)$. Recall that $p_{\log_{\delta}(a)}=\operatorname{tp}(a/L)$ is, by assumption, weakly orthogonal to $\mathcal{C}$. As $b\in\operatorname{acl}(aL)$, so is $\operatorname{tp}(a,b/L)$, which implies $L(b)(a)\cap\mathcal{C}=L\cap\mathcal{C}$, and thus $\frac{a^{n-i}}{c_{i}}\in L$, so $a^{n-i}\in L(b)$. If all $c_{i}$ are zero, then $a^{n}=0\in L(b)$. In any case, there is $k>0$ such that $a^{k}\in L(b)$.
Thus there are $f,g\in L[x]$ such that $a^{k}=\frac{f(b)}{g(b)}$. Taking derivatives, we obtain
$a^{k}=\frac{\delta(f(b))g(b)-f(b)\delta(g(g))}{k\log_{\delta}(a)g(b)^{2}}$
and thus, by equating our two expressions for $a^{k}$:
$\delta(f(b))g(b)-(k\log_{\delta}(a)g(b)+\delta(g(b)))f(b)=0\text{ .}$
Write $f(x)=\sum\limits_{i=0}^{n}e_{i}x^{i}$ and $g(x)=\sum\limits_{j=0}^{m}d_{j}x^{j}$ for some $e_{i},d_{j}\in L$ (with $e_{n}\neq 0$ and $d_{m}\neq 0$), we compute:
$\delta(f(b))$ $=\sum\limits_{i=0}^{n}\left(\delta(e_{i})+ie_{i}\log_{\delta}(b)\right)b^{i}$
$\delta(g(b))$ $=\sum\limits_{j=0}^{m}\left(\delta(d_{j})+jd_{j}\log_{\delta}(b)\right)b^{j}$
In particular, since $\log_{\delta}(b)\in L$, these expressions are given by polynomials in $L(x)$ applied to $b$. Therefore the equation (4.1):
$\delta(f(b))g(b)-(k\log_{\delta}(a)g(b)+\delta(g(b)))f(b)=0$
is a polynomial in $b$ over $L$. Since $\operatorname{RM}(b/L)=1$, all coefficients must be zero. In particular, examining the dominant coefficient, and using that $c_{n}\neq 0$ and $d_{m}\neq 0$, we get:
$\log_{\delta}(e_{n})+n\log_{\delta}(b)-k\log_{\delta}(a)-\log_{\delta}(d_{m})-m\log_{\delta}(b)=0$
which simplifies into:
$\log_{\delta}\left(a^{k}\right)=\log_{\delta}\left(\frac{e_{n}}{d_{m}}b^{n-m}\right)$
and implies that there is $\mu\in\mathcal{C}$ such that:
$a^{k}$ $=\mu\frac{e_{n}}{d_{m}}b^{n-m}$
$=\frac{e_{n}}{d_{m}}\left(\mu b^{n-m}\right)$
Note that $\log_{\delta}\left(\mu b^{n-m}\right)=(n-m)\log_{\delta}(b)\in F$. So setting $w_{1}=\frac{e_{n}}{d_{m}}$ and $w_{2}=\mu b^{n-m}$ gives us what we want.
As a corollary, we obtain a refinement of Theorem 4.4 of [10]:
@@
There exist a differential field $F$ and a minimal type $p\in S(F)$ internal to the constants, with a binding group of dimension $3$, such that $\log_{\delta}^{-1}(p)$ is $\mathcal{C}$-internal but $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ does not almost-splits. In particular the map $\log_{\delta}$ is uniformly internal but does not almost split.
@@
The first part is an immediate consequence of Theorem 4.4 of [10] and Theorem 4.5, the second part of [5, Proposition 3.16] showing the equivalence of internality of $\log_{\delta}^{-1}(p)$ and uniform internality of $f$. ∎
In other words, if a type is 2-analysable over $\mathcal{C}$, and also almost $\mathcal{C}$-internal, it need not split into a Morley product of almost $\mathcal{C}$-internal types.
The following lemma will be key in applications:
@@
Let $F$ be an algebraically closed field, let $q\in S(F)$ be an almost $\mathcal{C}$-internal type and $f:q\rightarrow f(q)=p$ be a fibration, with $p$ being $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal. Suppose that for some (any) $a\models q$:
  • $q_{f(a)}$ is $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal,
  • $\mathrm{Aut}_{f(a)F}(q_{f(a)}/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$.
Then there exists a type $q^{\prime}$, internal to $\mathcal{C}$, and a fibration $\log_{\delta}:q^{\prime}\rightarrow\log_{\delta}(q^{\prime})$ with $\log_{\delta}(q^{\prime})\subset\operatorname{dcl}(p(\mathcal{U})F)$. Moreover if $\log_{\delta}:q^{\prime}\rightarrow\log_{\delta}(q^{\prime})$ almost splits then $f:q\rightarrow f(q)$ almost splits.
Proof.
By Fact 3.7, there is a finite-to-one definable map $\pi:q\rightarrow\pi(q)$ such that $\pi(q)$ is $\mathcal{C}$-internal. In particular, for any $a\models q$, it restricts to a finite-to-one map from $q_{f(a)}$ to $\operatorname{tp}(\pi(a)/f(a)F)$. By Lemma 2.10, there is a finite-to-one definable morphism $\mathrm{Aut}_{f(a)F}(q_{f(a)}/\mathcal{C})\rightarrow\mathrm{Aut}_{f(a)F}(\operatorname{tp}(\pi(a)/f(a)F)/\mathcal{C})$. As the former group is definably isomorphic to $G_{m}(\mathcal{C})$, so is the latter.
Fix some $a\models q$. Since $q_{f(a)}$ is weakly $\mathcal{C}$-orthogonal, so is $\operatorname{tp}(\pi(a)/f(a)F)$. Moreover, as $p=f(q)$ is weakly orthogonal to $\mathcal{C}$, we have that $\mathcal{C}\cap F\langle f(a)\rangle=\mathcal{C}\cap F$, which is algebraically closed. By [10, Lemma 4.1], there is $b\in F\langle f(a),\pi(a)\rangle\setminus F\langle f(a)\rangle^{\mathrm{alg}}$ such that $\log_{\delta}(b)\in F\langle f(a)\rangle$. We let $q^{\prime}=\operatorname{tp}(b/F)$ and obtain a fibration $\log_{\delta}:q^{\prime}\rightarrow\log_{\delta}(q^{\prime})$ with $\mathcal{C}$-internal strongly minimal fibers. Finally as $b\in\operatorname{dcl}(f(a),\pi(a))$, we see that $q^{\prime}$ is $\mathcal{C}$-internal.
[…]
Claim.
We have $U(c/F)=1$.
Proof of claim.
[…]
$=U(c\log_{\delta}(b)/F)-U(\log_{\delta}(b)/F)$
$=U(b/F)-U(b/F)+1$
$=1$
[…]
By construction $b\in\operatorname{dcl}(aF)$, and $c\in\operatorname{acl}(bF)$, so $(c,f(a))\in\operatorname{acl}(aF)$. Note that $U(c,f(a)/F)=U(c/F)+U(f(a)/F)=1+U(f(a)/F)$. Since $\mathrm{Aut}_{f(a)F}(q_{f(a)}/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$ and $q_{f(a)}$ is weakly $\mathcal{C}$-orthogonal, we see that $U(q_{f(a)})=1$, which implies that $U(f(a)/F)=U(a/F)-1$, and combining this with the previous equality we get $U(c,f(a)/F)=U(a/F)$. As $(c,f(a))\in\operatorname{acl}(aF)$, this implies that $a$ and $(c,f(a))$ are interalgebraic over $F$, and $f:q\rightarrow f(q)$ splits.
Finally, we prove a criteria for almost splitting of generic types of differential equations of arbitrary high order. Recall that if $(F,\delta)$ is any differential field, then the field $F(x_{0},\cdots,x_{m})$ can be equipped with the derivations $\frac{\partial}{\partial x_{i}}$ with respect to $x_{i}$, as well as the derivation $\delta^{F}$, which treats all $x_{i}$ as constants and equals $\delta$ on $F$. If $P\in F[x_{0},\cdots,x_{m}]$, then $\delta^{F}(P)=P^{\delta}$, i.e. the polynomial obtained by differentiating the coefficients of $P$. Using this notation, we can compute, for any tuple $(a_{0},\cdots,a_{m-1})\in\mathcal{U}^{m}$ and $h\in F(x_{0},\cdots,x_{m-1})$, that $\delta(h(\overline{a}))=\sum\limits_{i=0}^{m-1}\frac{\partial h}{\partial x_{i}}\delta(a_{i})+\delta^{F}(h)(\overline{a})$. We obtain the following:
Lemma 4.8.
Let $F$ be an algebraically closed differential field, a rational function $f\in F(x_{0},\cdots,x_{m-1})$ and $q$ the generic type of $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$. Let $p:=\log_{\delta}^{-1}(q)$, then $\log_{\delta}:p\rightarrow q$ is almost split if and only if there are a non-zero $h\in F(x_{0},\cdots,x_{m-1})$, some $e\in F$ and some integer $k\neq 0$ such that
@@
For the converse, suppose that $\log_{\delta}:p\rightarrow q$ is almost split. By Theorem 4.5, there are $u\models p$, some integer $k\neq 0$ and $w_{1},w_{2}$ such that:
@@ 4.2. Splitting when the binding group is nilpotent
In this subsection, we prove that for any $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal type $p$ over any algebraically closed field $F$, if $p$ has nilpotent binding group, then $\log_{\delta}^{-1}(p)$ is almost $\mathcal{C}$-internal if and only if $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ splits. This can be seen as a partial generalization of Theorem A of [10].
@@
Let $F$ be an algebraically closed differential field and let $p\in S_{1}(F)$ be a $\mathcal{C}$-internal, weakly $\mathcal{C}$-orthogonal type with binding group definably isomorphic to the $\mathcal{C}$-points of a nilpotent linear algebraic group. Then $\log_{\delta}^{-1}(p)$ is almost $\mathcal{C}$-internal if and only if $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ almost splits.
@@
That almost splitting implies almost internality is immediate. So assume that $q=\log_{\delta}^{-1}(p)$ is almost $\mathcal{C}$-internal. To use binding groups, we need to reduce to the case where $q$ is $\mathcal{C}$-internal.
@@
First note that if $q_{\log_{\delta}(a)}$ is not weakly $\mathcal{C}$-orthogonal, by Lemma 4.4 $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ is almost split. So we can assume that $q_{\log_{\delta}(a)}$ is weakly $\mathcal{C}$-orthogonal. By Lemma 4.7, there is another type $q^{\prime}$, internal to $\mathcal{C}$, and a fibration $\log_{\delta}:q^{\prime}\rightarrow\log_{\delta}(q^{\prime})$ with $\log_{\delta}(q^{\prime})\subset\operatorname{dcl}(p(\mathcal{U})F)$. This last part implies that the binding group of $\log_{\delta}(q^{\prime})$ is also definably isomorphic to the $\mathcal{C}$-points of a nilpotent linear algebraic group. Moreover if $\log_{\delta}:q^{\prime}\rightarrow\log_{\delta}(q^{\prime})$ almost splits, so does $\log_{\delta}:q\rightarrow\log_{\delta}(q)$. Therefore, replacing $q$ by $q^{\prime}$, we may assume that $q$ is $\mathcal{C}$-internal.
Hence we get a short exact sequence:
$1\rightarrow K\rightarrow\mathrm{Aut}_{F}(q/\mathcal{C})\xrightarrow{\widetilde{\log_{\delta}}}\mathrm{Aut}_{F}(p/\mathcal{C})\rightarrow 1$
By [6, Lemma 2.7] $K$ is definably isomorphic to $\mathrm{Aut}_{F}(\operatorname{tp}(\bar{a}/F\log_{\delta}(\bar{a}))/\mathcal{C})$, where $\bar{a}=a_{1},\cdots,a_{n}$ is Morley sequence in $q$ such that $\bar{a}$ is a fundamental system of realizations of $q$, and $\log_{\delta}(\bar{a})$ is a fundamental system of realizations for $p$. As $\mathrm{Aut}_{F\log_{\delta}(a)}(\operatorname{tp}(a/F\log_{\delta}(a))/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$ for any $a\models q$, this implies that $K$ is definably isomorphic to a subgroup of $G_{m}(\mathcal{C})^{n}$. As $F$ is algebraically closed, it is connected, so in fact isomorphic to a torus. Thus $\mathrm{Aut}_{F}(q/\mathcal{C})$ is definably isomorphic to the extension of a nilpotent linear algebraic group by a torus, thus by Lemma 2.18 must be a isomorphic to the $\mathcal{C}$-point of a nilpotent linear algebraic group.
Moreover, remark that the maximal tori of $\mathrm{Aut}_{F}(p/\mathcal{C})$ and $\mathrm{Aut}_{F}(q/\mathcal{C})$ are $F$-definably isomorphic to the $\mathcal{C}$-points of tori. Indeed, because both groups are definably nilpotent, their maximal tori are central. We can modify the proof of [6, Lemma 2.2] to show that they must then be $F$-definably isomorphic to the $\mathcal{C}$-points of tori.
By Lemma 2.18, the short exact sequence of binding groups splits $F$-definably, let $s$ be the section. We will use this to show that the groupoid $\mathrm{Gd}_{F}(q,\log_{\delta}/\mathcal{C})$ is retractable.
We know that $G=\mathrm{Aut}_{F}(p/\mathcal{C})=G_{u}\times T$, where $T$ is its maximal torus. Consider the projections $\pi_{1}$ and $\pi_{2}$ on $G_{u}$ and $T$.
Claim.
Proof.
By Lemma 2.9 applied to the group $\pi_{1}(H)\times\pi_{2}(H)$, there is a definable isomorphism between $\pi_{1}(H)/N_{1}$ and $\pi_{2}(H)/N_{2}$, where $N_{1}=\ker(\pi_{2}|_{H})$ and $N_{2}=\ker(\pi_{1}|_{H})$ are normal definable subgroups of $H$, which we can identify with subgroups of $\pi_{1}(H)$ and $\pi_{2}(H)$. But the group $\pi_{1}(H)/N_{1}$ is definably unipotent and the group $\pi_{2}(H)/N_{2}$ is a definable torus, so there cannot be any definable isomorphism between them, unless they are both trivial. Therefore $\pi_{1}(H)=N_{1}$ and $\pi_{2}(H)=N_{2}$, which is equivalent to $H=\pi_{1}(H)\times\pi_{2}(H)$. ∎
The action of $G$ on $p$ is definably isomorphic to its action on $G/H$, where $H<G$ is a definable subgroup. By the claim, we know that $H=\pi_{1}(H)\times\pi_{2}(H)$, and thus the action is definably isomorphic to the action of $G=G_{u}\times T$ on $G_{u}/\pi_{1}(H)\times T/\pi_{2}(H)$. In particular for any $a\models p$, we have $G_{u}a\cap Ta=\{a\}$. Because $G_{u}$ is unipotent, by Lemma 3.17, its action on $p$ satisfies (normal-$\bigstar$). Because $T$ is abelian, its action also satisfies (normal-$\bigstar$). Therefore the action of $G$ satisfies (normal-$\bigstar$). By Lemma 3.16, the map $\log_{\delta}:q\rightarrow p$ splits.
@@
This is an immediate consequence of Lemma 4.8 and Theorem 4.9.
@@
Let $v_{0}$ be the valuation in $F(x_{0},\cdots,x_{m-1})$ with respect to $x_{0}$. The coefficients of $x_{0}^{-v_{0}(h)+1}$ must be equal on both sides of the equality. We can write $h=\widetilde{h}x_{0}^{-v_{0}(h)}+g(x_{0},\cdots,x_{m-1})$ with $\widetilde{h}\in F(x_{1},\cdots,x_{m-1})$ and $v_{0}(g)>v_{0}(h)$. We have:
@@
If $F$ is a field of constants, then this would apply as long as the generic type of $\delta^{m}(x)=L(\delta^{m-1}(x),\cdots,\delta(x),x)$ is weakly $\mathcal{C}$-orthogonal, as its binding group is always abelian in that case (see the proof of [1, Theorem 3.9] for a justification). Note that this is not always the case: for example, the generic type of $\delta^{m}(x)=0$ is never weakly $\mathcal{C}$-orthogonal as long as $m>1$.

### v3

@@ 4.1. Preliminaries on logarithmic-differential pullbacks
We have that $a$ and $b$ are non-zero, algebraically dependent over $L$, and both $\log_{\delta}(a)\in L$ and $\log_{\delta}(b)\in L$. By the Kolchin-Ostrowski Theorem (see [16, top of page 1156]), we can conclude that there is some $w_{1}\in L$, some $k,l\in\mathbb{Z}$ not both zero, such that $a^{k}b^{l}=w_{1}$. Note that $l\neq 0$, as otherwise $a\in L^{\mathrm{alg}}$, a contradiction. Therefore $k\neq 0$ and we get the conclusion by picking $w_{2}=b^{-l}$.
Over a field of constants, we obtain the equivalence in general :
@@
Let $q\in S_{1}(F)$ be a type over any algebraically closed differential field of constants. Let $p:=\log_{\delta}^{-1}(q)$. Then $\log_{\delta}:p\to q$ is almost split if and only if there is an integer $k\neq 0$ such that for some $a\models p$, there are $w_{1},w_{2}$ such that:
  • $a^{k}=w_{1}w_{2}$
  • $w_{1}\in\operatorname{dcl}(F,\log_{\delta}a)$
  • $\log_{\delta}(w_{2})\in F$.
@@
As before, we only have to prove that almost splitting implies product-splitting. In the previous proof, the only time we used the extra assumption was to show that the binding group of $s$ was definably isomorphic to $G_{m}(\mathcal{C})$. Keeping the notation of that proof, we now show how to obtain this if $F$ is a field of constants.
By Fact 2.5, the binding group $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$ is a definable subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$. Moreover, we still know that $\mathrm{Aut}_{F}(s/\mathcal{C})$ acts transitively and faitfhfully on the strongly minimal type $s$. Over the constants, the only possible binding groups in Fact 4.2 are $G_{a},G_{m}$ and an elliptic curve (see the proof of [3, Theorem 3.9]), and therefore the only possible case is $\mathrm{Aut}_{F}(s/\mathcal{C})$ being definably isomorphic to $G_{m}(\mathcal{C})$. ∎
We also always obtain the equivalence if the base type is orthogonal to the constants:
Corollary 4.7.
Let $q\in S_{1}(F)$ be a type over any algebraically closed differential field, and assume that $q$ is orthogonal to the constants. Let $p:=\log_{\delta}^{-1}(q)$. Then $\log_{\delta}:p\to q$ is almost split if and only if there is an integer $k\neq 0$ such that for some $a\models p$, there are $w_{1},w_{2}$ such that:
  • $a^{k}=w_{1}w_{2}$
  • $w_{1}\in\operatorname{dcl}(F,\log_{\delta}a)$
  • $\log_{\delta}(w_{2})\in F$.
Proof.
Again, we keep the notation of the proof of Theorem 4.5. By that theorem, we may assume that $p(\mathcal{U})\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$, which implies that $s(\mathcal{U})\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$.
As $q=\log_{\delta}(p)$ is orthogonal to the constants, we have that $\mathrm{Aut}_{F}(s/\mathcal{C},\log_{\delta}(p))=\mathrm{Aut}_{F}(s/\mathcal{C})$. The former group is, by assumption, finite. Thus $s$ is $\mathcal{C}$-algebraic. This implies, in particular, that for some (any) $a\models p$, the fiber $p_{\log_{\delta}(a)}$ is not weakly $\mathcal{C}$-orthogonal. By Lemma 4.4, we get product-splitting. ∎
Finally, we prove a criteria for product-splitting of generic types of differential equations of arbitrary high order. More specifically, if $F$ is some algebraically closed differential subfield of $\mathcal{U}$, we will be interested in definable sets given by solution sets of equations of the form $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$, with $f\in F(x_{0},\cdots,x_{m-1})$. Note that we can write $f=\frac{P}{Q}$ with $P,Q\in F[x_{0},\cdots,x_{m-1}]$ without common factors, and that our definable set is then given by solutions of the equation $\delta^{m}(y)Q(y,\delta(y),\cdots,\delta^{m-1}(y))-P(y,\delta(y),\cdots,\delta^{m-1}(y))=0$. As the polynomial $x_{m}P(x_{0},\cdots,x_{m-1})-Q(x_{0},\cdots,x_{m-1})\in F[x_{0},\cdots,x_{m}]$ is irreducible, this definable set has a unique generic type.
Recall that if $(F,\delta)$ is any differential field, then the field $F(x_{0},\cdots,x_{m})$ can be equipped with the derivations $\frac{\partial}{\partial x_{i}}$ with respect to $x_{i}$, as well as the derivation $\delta^{F}$, which treats all $x_{i}$ as constants and equals $\delta$ on $F$. If $P\in F[x_{0},\cdots,x_{m}]$, then $\delta^{F}(P)=P^{\delta}$, i.e. the polynomial obtained by differentiating the coefficients of $P$. Using this notation, we can compute, for any tuple $(a_{0},\cdots,a_{m-1})\in\mathcal{U}^{m}$ and $h\in F(x_{0},\cdots,x_{m-1})$, that $\delta(h(\overline{a}))=\sum\limits_{i=0}^{m-1}\frac{\partial h}{\partial x_{i}}\delta(a_{i})+\delta^{F}(h)(\overline{a})$. We obtain the following:
@@
Let $F$ be an algebraically closed differential field, a rational function $f\in F(x_{0},\cdots,x_{m-1})$ and $q$ the generic type of $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$. Let $p:=\log_{\delta}^{-1}(q)$, then $\log_{\delta}:p\rightarrow q$ is product-split if and only if there are a non-zero $h\in F(x_{0},\cdots,x_{m-1})$, some $e\in F$ and some integer $k\neq 0$ such that
@@
For the converse, suppose that $\log_{\delta}:p\rightarrow q$ is product-split. There are $u\models p$, some integer $k\neq 0$ and $w_{1},w_{2}$ such that:
@@
Under the additional assumption of Theorem 4.5, we get a criteria for almost splitting:
Corollary 4.9.
Let $F$ be an algebraically closed differential field, consider some $f\in F(x_{0},\cdots,x_{m-1})$ and $q$ the generic type of $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$. Let $p:=\log_{\delta}^{-1}(q)$ and assume that $p(\mathcal{U})\not\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$. Then $\log_{\delta}:p\rightarrow q$ is almost split if and only if there are a non-zero $h\in F(x_{0},\cdots,x_{m-1})$, some $e\in F$ and some integer $k\neq 0$ such that
Proof.
Immediate consequence of Theorem 4.5 and Lemma 4.11. ∎
@@ 4.2. Splitting when the binding group is nilpotent
In this subsection, we prove that for any $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal type $q$ over any algebraically closed field $F$, if $q$ has nilpotent binding group, then $\log_{\delta}^{-1}(q)$ is almost $\mathcal{C}$-internal if and only if $\log_{\delta}:\log_{\delta}^{-1}(q)\rightarrow q$ splits. This can be seen as a partial generalization of Theorem A of [12].
We will sometimes return to the notation of Subsection 2.2: for example, by a definable nilpotent linear group, we mean a definable group definably isomorphic to the $\mathcal{C}$-points of a nilpotent linear algebraic group.
@@
Let $F$ be an algebraically closed differential field and let $q\in S_{1}(F)$ be a $\mathcal{C}$-internal, weakly $\mathcal{C}$-orthogonal type with binding group definably isomorphic to the $\mathcal{C}$-points of a nilpotent linear algebraic group. Then the following are equivalent:
  1. (1)
    $\log_{\delta}^{-1}(q)$ is almost $\mathcal{C}$-internal,
  2. (2)
    $\log_{\delta}:\log_{\delta}^{-1}(q)\rightarrow q$ product-splits,
  3. (3)
    $\log_{\delta}:\log_{\delta}^{-1}(q)\rightarrow q$ almost splits.
@@
Assume that $p=\log_{\delta}^{-1}(q)$ is almost $\mathcal{C}$-internal. We will start by showing that this implies almost splitting of $\log_{\delta}$. To use binding groups, we need to reduce to the case where $p$ is $\mathcal{C}$-internal.
First note that if $p_{\log_{\delta}(a)}$ is not weakly $\mathcal{C}$-orthogonal, by Lemma 4.4 $\log_{\delta}:\log_{\delta}^{-1}(q)\rightarrow q$ product-splits. So we now assume that $p_{\log_{\delta}(a)}$ is weakly $\mathcal{C}$-orthogonal. As $q$ is also weakly $\mathcal{C}$-orthogonal, we see that $p$ is weakly $\mathcal{C}$-orthogonal.
@@
By Lemma 4.3, there is another type $p^{\prime}$, internal to $\mathcal{C}$, and a fibration $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ with $\log_{\delta}(p^{\prime})\subset\operatorname{dcl}(q(\mathcal{U})F)$, with the property that if $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ almost splits then $\log_{\delta}:p\rightarrow\log_{\delta}(p)$ almost splits. Because $p$ is weakly $\mathcal{C}$-orthogonal, the proof of Lemma 4.3 shows that $p^{\prime}$ also is.
We see that $\log_{\delta}(p^{\prime})$ is weakly $\mathcal{C}$-orthogonal, and its binding group also is definably isomorphic to the $\mathcal{C}$-points of a nilpotent linear algebraic group. Moreover if $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ almost splits, so does $\log_{\delta}:p\rightarrow\log_{\delta}(p)$. Therefore, replacing $p$ by $p^{\prime}$, we may assume that $p$ is $\mathcal{C}$-internal. By Theorem 3.18, the fibration $\log_{\delta}:p\rightarrow q$ is almost split.
We still need to obtain product-splitting. By Theorem 4.5, we may assume that $p(\mathcal{U})\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$. Using similar ideas as in the proof of Corollary 2.13, we see that this implies that $\mathrm{Aut}_{F}(p/\mathcal{C})$ is finite. This group is the kernel of $\log_{\delta}:p\rightarrow\log_{\delta}(p)$. By Lemma 3.17, we obtain that the fibers of $f$ are not weakly $\mathcal{C}$-orthogonal, which is a contradiction, as we assumed that they were.
@@
This is an immediate consequence of Lemma 4.8 and Theorem 4.10. ∎
@@
Let $v_{0}$ be the valuation in $F(x_{0},\cdots,x_{m-1})$ with respect to $x_{0}$. We can write $h=\widetilde{h}x_{0}^{-v_{0}(h)}+g(x_{0},\cdots,x_{m-1})$ with $\widetilde{h}\in F(x_{1},\cdots,x_{m-1})$ and $v_{0}(g)>v_{0}(h)$. We have:
@@
If $F$ is a field of constants, then this would apply as long as the generic type of $\delta^{m}(x)=L(\delta^{m-1}(x),\cdots,\delta(x),x)$ is weakly $\mathcal{C}$-orthogonal, as its binding group is always abelian in that case (see the proof of [3, Theorem 3.9] for a justification). Note that this is not always the case: for example, the generic type of $\delta^{m}(x)=0$ is never weakly $\mathcal{C}$-orthogonal as long as $m>1$.